Example A: 2n + 3 = 13
Example B: n + 3 = 13
Which example is a one-step equation?
How is solving a two-step equation like solving one-step equation? How is solving them different?

Answer by ewatrrr(24785)


Hi
Example A: 2n + 3 = 13 |two-step
Example B: n + 3 = 13 |one-step (subtracting 3 from both sides) n = 10
Likeness: both involve isolating the variable on one side of the equation
Difference: A second step is necessary in the two-step to accomplish this.
for example:
2n + 3 = 13
2n = 10 |Step one, subtracting 3 from both sides
n = 5 |Step two, dividing both sides by 2